# Upper and Lower Bounds for the Linear Ordering Principle

### v1

@@ 1.3.2 A New Upper Bound for $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$
Containment.
Our second main result is a new upper bound on $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. In particular, we prove that $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{P}}^{\mathbf{{prSBP}}}$. The best known upper bound prior to our result followed from [KP24, CR11]: $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{P}}^{\mathbf{{prAM}}}$.
@@
In summary, our two new inclusions yield the following picture:
$\mathbf{{MA}}$$\mathbf{{SBP}}$$\mathbf{{AM}}$$\mathbf{{P}}^{\mathbf{{prMA}}}$$\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$$\mathbf{{P}}^{\mathbf{{prSBP}}}$$\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$$\mathbf{{P}}^{\mathbf{{prAM}}}$heorem 1Theorem 4
Implications for “hard” functions.
In [KP24], Korten and Pitassi have shown that $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle E}}}}}$ — the exponential version of $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ — contains a language of circuit complexity $2^{n}/n$. By translation, our upper bound scales as $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle E}}}}}\subseteq\mathbf{{E}}^{\mathbf{{prSBP}}}$. As a corollary we obtain new circuit lower bounds for $\mathbf{{E}}^{\mathbf{{prSBP}}}$. To the best of our knowledge, the strongest previously established bound for this class was a “half-exponential” bound that followed from the bound on $\mathbf{{MAEXP}}$ [BFT98, MVW99].
@@ 1.4 Our Techniques
1.4.1 Range Avoidance
In [KP24], Korten and Pitassi show that $\mathbf{{MA}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. The inclusion $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ essentially follows their argument with the additional observation that since $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ is a syntactic class, not only it contains $\mathbf{{MA}}$ (as was shown) but it is also consistent with $\mathbf{{prMA}}$. Thus one can first construct a pseudorandom generator using an $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ oracle [Kor21, KP24] and then leverage it to fully derandomize the $\mathbf{{prMA}}$ oracle not just in $\mathbf{{prNP}}$, but actually in $\mathbf{{NP}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. Therefore, $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{P}}^{\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}=\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$.
1.4.2 Input-Oblivious Symmetric Alternation
A Karp–Lipton–style collapse to $\mathbf{{P}}^{\mathbf{{prOMA}}}$ follows from [CR11] by combining several previously known techniques. However, is this collapse stronger than the known collapse to $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [CR06]? The inclusion $\mathbf{{prOMA}}\subseteq\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ can be transferred from a somewhat similar statement that was proven in [CR06]; however, in order to prove $\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ we need also the inclusion $\mathbf{{P}}^{\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, which seems novel. The main idea is that the two provers corresponding to the oracle give their input-oblivious witnesses prior to the whole computation, and the verification algorithm performs a cross-check not only between the witnesses of different provers but also between the witnesses of the same prover, which allows us to simulate all oracle queries to $\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ in a single $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ algorithm. Indeed, our approach is made possible by the input-oblivious nature of the computational model: while the oracle queries may be adaptive and not known in advance (due to potential queries outside of the promise set), the witnesses are universal for the whole computation and nothing else is required.
@@ 1.4.3 Approximate Counting and the Order Rank Approximation
Approximating the rank w.r.t. a linear order.
The rank of an element $\alpha$ of a linearly ordered set $U$ is the number of elements in this set that are strictly less than $\alpha$ (in particular, $\alpha$ is the minimum if and only if $\mathrm{rank}(\alpha)=0$). We can extend this definition to non-empty subsets $S\subseteq U$, where $\mathrm{rank}(S)$ is the average rank of elements in $S$.
@@
By averaging argument, $\min{\left\{\mathrm{rank}(S_{0}),\mathrm{rank}(S_{1})\right\}}\leq\mathrm{rank}(S)$. We then take $S$ to be the subset ($S_{0}$ or $S_{1}$) with the smaller rank and continue to the next value of $i$. That is, we fix the bits of the elements of $S$ one coordinate at a time. Therefore, once $i=n$, our “final” set $S$ contains exactly one element $\beta$ and thus at that point $\mathrm{rank}(S)=\mathrm{rank}(\beta)$. On the other hand, as the rank of the “initial” $S$ was $\mathrm{rank}(\alpha)/2$ and the overall rank could only decrease, we obtain that $\mathrm{rank}(\beta)\leq\mathrm{rank}(\alpha)/2$. We can then invoke the same procedure this time with $\beta$ as its input. As the there are $2^{n}$ elements in $U$, this process will converge to the set’s minimum after invoking the procedure at most $n$ times, given any initial element.
The algorithm described above requires computing (or at least comparing) the average ranks of two sets. Our analysis demonstrates that a procedure for approximate comparison, developed before, is sufficient for the implementation of this idea (though the factor at each step will be a little bit less than $2$).

### v2

@@ 1.3.2 A New Upper Bound for $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$
Another important result of this work is a new upper bound on $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$: we prove that $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{P}}^{\mathbf{{prSBP}}}$. The best known upper bound prior to our result followed from [KP24, CR11]: $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{P}}^{\mathbf{{prAM}}}$.
@@
In summary, our two new inclusions (Theorems 1 and 3) yield the “normal” (non-input-oblivious) part of Figure 1.
[…]
Figure 1: Containments of classes based on Merlin–Arthur protocols and on symmetric alternation.
Switching to the linear-exponential regime, in [KP24], Korten and Pitassi have shown that $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle E}}}}}$ — the exponential version of $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ — contains a language of circuit complexity $2^{n}/n$. By translation, our upper bound scales as $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle E}}}}}\subseteq\mathbf{{E}}^{\mathbf{{prSBP}}}$  [^4] As is turns out, this class is also equal $\mathbf{{E}}^{\mathbf{{prBPP_{path}}}}$. See Section 1.4.1 for more details.. As a corollary we obtain a new circuit lower bound for the class $\mathbf{{E}}^{\mathbf{{prSBP}}}$. To the best of our knowledge, the strongest previously established bound for this class was “half-exponential” that followed from the bound on $\mathbf{{MAEXP}}$ [MVW99].
@@
1.3.3 Aggregation of $\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ queries and the Strongest Input-Oblivious Karp–Lipton Collapse
Theorem 1 shows that $\mathbf{{P}}^{\mathbf{{prMA}}}$ is currently the smallest non-input-oblivious class for which a Karp–Lipton–style collapse is known. On the other hand, such a collapse was also shown for $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [CR06], which is input-oblivious. However, since the precise relationship between $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ and $\mathbf{{P}}^{\mathbf{{prMA}}}$ remains unknown, one may ask: what is the strongest Karp–Lipton–style collapse? Our next result assists in navigating this question.
Theorem 5.
We note that the “non-promise” version of this inclusion, i.e. $\mathbf{{P}}^{\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, was already established in [CR06]. However, this result does not carry over to the promise case. A similar phenomenon arises in the non-input-oblivious analogue of this question: while we know that $\mathbf{{P}}^{\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [CR06], it still remains open whether $\mathbf{{P}}^{\mathbf{{prS_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$.
With this tool in hand, we can identify and show the strongest Karp–Lipton–style collapse that currently known. Namely, the collapse can be extended to $\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{P}}^{\mathbf{{prMA}}}$, where $\mathbf{{prOMA}}$ is the input-oblivious version of $\mathbf{{prMA}}$ and therefore is contained in $\mathbf{{prMA}}$. At the same time, Theorem 5 allows us to show that $\mathbf{{P}}^{\mathbf{{prOMA}}}$ is also contained in $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, thus making $\mathbf{{P}}^{\mathbf{{prOMA}}}$ smaller than both $\mathbf{{P}}^{\mathbf{{prMA}}}$ and $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$!
Theorem 5 also allows us to refine the chain of containments between $\mathbf{{P}}^{\mathbf{{prOMA}}}$ and $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ by introducing a newly defined class, $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, an input-oblivious analogue of $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ (see Figure 1). Note that this chain proceeds via $\mathbf{{P}}^{\mathbf{{ONP}},\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$ rather than $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ itself, since $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ is not guaranteed either to share the convenient closure properties of $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ or to contain $\mathbf{{ONP}}$. Still we prove that
$\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{P}}^{\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}},\mathbf{{prONP}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\cap\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}.$
That is, $\mathbf{{P}}^{\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}},\mathbf{{ONP}}}$ can serve as an input-oblivious analogue of $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ ($\mathbf{{P}}^{\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}},\mathbf{{NP}}}$ [KP24]).
@@ 1.4 Our Techniques
1.4.1 Approximate Counting, Set Size Estimation and $\mathbf{{SBP}}$
Computing the number of accepting paths of a given non-deterministic Turing machine is a fundamental problem captured by the “counting” class $\mathbf{\#\mathbf{{P}}}$. Yet, this class appears to be too powerful since, by Toda’s Theorem [Tod91], even a single query to it suffices to decide any language in the polynomial hierarchy, $\mathbf{{PH}}\subseteq\mathbf{{P}}^{\mathbf{\#\mathbf{{P}}[1]}}$! Given that, it is natural to explore approximations. To this end, one can consider the problem of Approximate Counting (ApproxCount for short) which refers to the task of approximating the number of accepting paths (within a constant factor). Equivalently, this problem can be framed as approximating the size of a set $S$ represented as the set of satisfying assignments of a Boolean circuit $C$. Previously, it was shown that this task could be carried out by a randomized algorithm using an $\mathbf{{NP}}$ oracle ($\mathbf{{FBPP}}^{\mathbf{{NP}}}$) [Sto85, JVV86] and by a deterministic algorithm using a $\mathbf{{prAM}}$ oracle ($\mathbf{{FP}}^{\mathbf{{prAM}}}$) [Sip83, GS86]. Shaltiel and Umans [SU06] show how to accomplish this task in $\mathbf{{FP}}^{\mathbf{{NP}}}$, yet under a derandomization assumption. We note that all of these algorithms can be implemented using parallel (i.e. non-adaptive) oracle queries. That is, in $\mathbf{{FBPP}}^{\mathbf{{NP}}}_{\parallel},\mathbf{{FP}}^{\mathbf{{prAM}}}_{\parallel}$ and $\mathbf{{FP}}^{\mathbf{{NP}}}_{\parallel}$, respectively. Approximate counting has also recently attracted considerable attention in the quantum literature [OS18, AKKT20, AR20, MAD25].
The decision version of the problem is to distinguish between two constant-factor estimates of the set size. For concreteness, consider the following problem called set-size estimation (or SSE for short): Given a set $S$ (via a Boolean circuit $C$) and an integer $m$ with the promise that either $\left|{S}\right|\geq m$ or $\left|{S}\right|\leq m/2$, our goal is to decide which case holds. Interestingly, this problem is complete for the class $\mathbf{{SBP}}$ [^6] Strictly speaking, the problem is complete for $\mathbf{{prSBP}}$, the corresponding class of promise problems, and is therefore hard for $\mathbf{{SBP}}$. introduced in [BGM06] by Böhler et al. as a relaxation of the class $\mathbf{{BPP}}$ to the case when the acceptance probability is not required to be bounded away from 0. This relaxation, as was shown in [BGM06], yields additional power: $\mathbf{{MA}}\subseteq\mathbf{{SBP}}\subseteq\mathbf{{AM}}$. The class $\mathbf{{SBP}}$, which stands for small bounded-error polynomial-time, thus sits strictly between the two fundamental classes based on Arthur–Merlin protocols, yet its definition is not based on these protocols. Moreover, $\mathbf{{SBP}}$ remains the only known natural class that lies between $\mathbf{{MA}}$ and $\mathbf{{AM}}$.
In terms of upper bounds, in a seminal paper [GS86], Goldwasser and Sipser have exhibited an Arthur–Merlin protocol not only for this problem, but also for the case when the set $S$ is represented by a non-deterministic circuit [^7] A non-deterministic circuit $C(x,w)$ accepts $x$ if there exists a witness $w$ for which C(x,w)=1C(x,w)=1.! This more general version of the problem (WSSE, where the set $S$ is given via a non-deterministic circuit), is complete for the class $\mathbf{{prAM}}$. (See Definition 3.1 for the formal definition of the problems; in fact, the factor-of-two gap in the estimates is arbitrary and can be replaced by any positive constant.) In fact, Goldwasser–Sipser’s protocol proves the containment both for languages ($\mathbf{{SBP}}\subseteq\mathbf{{AM}}$) and for promise problems ($\mathbf{{prSBP}}\subseteq\mathbf{{prAM}}$). At the same time, it is important to highlight the distinction between the two versions of the problem — i.e. for the “standard” (SSE) vs non-deterministic (WSSE) Boolean circuits — which appears to be (at the very least) non-trivial. Notably, the work of [BGM06] established an oracle separation between $\mathbf{{SBP}}$ and $\mathbf{{AM}}$.
On a similar note, by combining some of the previous techniques, we observe that ApproxCount can be carried out in $\mathbf{{FP}}^{\mathbf{{prSBP}}}$ rather than $\mathbf{{FP}}^{\mathbf{{prAM}}}$, and, in fact, even in $\mathbf{{FP}}^{\mathbf{{prSBP}}}_{\parallel}$. Given this observation, it is natural to study the computational power of $\mathbf{{P}}^{\textsc{ApproxCount}}$, that is, deterministic algorithms with oracle access to ApproxCount. Indeed, an immediate corollary of the above is that $\mathbf{{P}}^{\textsc{ApproxCount}}=\mathbf{{P}}^{\mathbf{{prSBP}}}$ and $\mathbf{{P}}^{\textsc{ApproxCount}}_{\parallel}=\mathbf{{P}}^{\mathbf{{prSBP}}}_{\parallel}$. At the same, O’Donnell and Say [OS18] previously showed that $\mathbf{{P}}^{\textsc{ApproxCount}}_{\parallel}=\mathbf{{BPP_{path}}}$, a complexity class defined earlier by Han et al. [HHT97]. One can think of $\mathbf{{BPP_{path}}}$ as a version of $\mathbf{{BPP}}$ in which different computational paths (of the same probability) may have different lengths. Incidentally, it was established in [BGM06] that $\mathbf{{BPP_{path}}}$ can be obtained from $\mathbf{{BPP}}$ via the so-called “postselection” and that $\mathbf{{SBP}}\subseteq\mathbf{{BPP_{path}}}$ (and resp. $\mathbf{{prSBP}}\subseteq\mathbf{{prBPP_{path}}}$). Putting all together, one arrives at the following three clusters of complexity classes associated with approximate counting:
[…]
where both inclusions are believed to be strict.
@@ 1.4.2 Approximate Counting and the Order Rank Approximation
Approximating the order rank w.r.t. a linear order.
The order rank of an element $\alpha$ of a linearly ordered set $U$ is the number of elements in this set that are strictly less than $\alpha$ (in particular, $\alpha$ is the minimum if and only if $\mathrm{rank}(\alpha)=0$). We can extend this definition to non-empty subsets $S\subseteq U$, where $\mathrm{rank}(S)$ is the average order rank of elements in $S$.
@@
By averaging argument, $\min{\left\{\mathrm{rank}(S_{0}),\mathrm{rank}(S_{1})\right\}}\leq\mathrm{rank}(S)$. We then take $S$ to be the subset ($S_{0}$ or $S_{1}$) with the smaller order rank and continue to the next value of $i$. That is, we fix the bits of the elements of $S$ one coordinate at a time. Therefore, once $i=n$, our “final” set $S$ contains exactly one element $\beta$ and thus at that point $\mathrm{rank}(S)=\mathrm{rank}(\beta)$. On the other hand, as the order rank of the “initial” $S$ was $\mathrm{rank}(\alpha)/2$ and the overall order rank could only decrease, we obtain that $\mathrm{rank}(\beta)\leq\mathrm{rank}(\alpha)/2$. We can then invoke the same procedure this time with $\beta$ as its input. As the there are $2^{n}$ elements in $U$, this process will converge to the set’s minimum after invoking the procedure at most $n$ times, given any initial element.
The algorithm described above requires computing (or at least comparing) the average order ranks of two sets. Our analysis demonstrates that a procedure for approximate comparison, developed before, is sufficient for the implementation of this idea (though the factor at each step will be a little bit less than $2$).
1.4.3 Derandomization in $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$
In [KP24], Korten and Pitassi show that $\mathbf{{MA}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. The inclusion $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ essentially follows their argument with the additional observation that since $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ is a syntactic class, not only it contains $\mathbf{{MA}}$ (as was shown) but it is also consistent with $\mathbf{{prMA}}$. Thus one can first construct a pseudorandom generator using an $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ oracle [Kor22, KP24] and then leverage it to fully derandomize the $\mathbf{{prMA}}$ oracle not just in $\mathbf{{prNP}}$, but actually in $\mathbf{{NP}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$! Therefore, $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{P}}^{\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}=\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$.
We also observe that since the pseudorandom generator depends only on the input length (and not the input itself), derandomization also helps settling the relations between input-oblivious classes to a certain extent. The main difference is that unlike their non-oblivious counterparts, they do not posses all the desired properties w.r.t. Turing closure and natural containments. However, Theorem 5 (its technique is described below) eventually helps us building the chain from $\mathbf{{P}}^{\mathbf{{prOMA}}}$ to $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ in two different ways: both directly and through derandomization and intermediate classes using $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$.
1.4.4 Input-Oblivious Symmetric Alternation
A Karp–Lipton–style collapse to $\mathbf{{P}}^{\mathbf{{prOMA}}}$ follows from [CR11] by combining several previously known techniques. However, is this collapse stronger than the known collapse to $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [CR06]? The inclusion $\mathbf{{prOMA}}\subseteq\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ can be transferred from a somewhat similar statement that was proven in [CR06]; however, in order to prove $\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ we need also the inclusion $\mathbf{{P}}^{\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, which seems novel. The main idea is that the two provers corresponding to the oracle give their input-oblivious certificates prior to the whole computation, and the verification algorithm performs a cross-check not only between the certificates of different provers but also between the certificates of the same prover, which allows us to simulate all oracle queries to $\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ in a single $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ algorithm. Indeed, our approach is made possible by the input-oblivious nature of the computational model: while the oracle queries may be adaptive and not known in advance (due to potential queries outside of the promise set), the certificates are universal for the whole computation and nothing else is required.